### v1

# Feature Manipulation for DDPM based Change Detection

Abstract
Change Detection is a classic task of computer vision that receives a bi-temporal image pair as input and separates the semantically changed and unchanged regions of it. The diffusion model is used in image synthesis and as a feature extractor and has been applied to various downstream tasks. Using this, a feature map is extracted from the pre-trained diffusion model from the large-scale data set, and changes are detected through the additional network. On the one hand, the current diffusion-based change detection approach focuses only on extracting a good feature map using the diffusion model. It obtains and uses differences without further adjustment to the created feature map. Our method focuses on manipulating the feature map extracted from the Diffusion Model to be more semantically useful, and for this, we propose two methods: Feature Attention and FDAF. Our model with Feature Attention achieved a state-of-the-art F1 score (90.18) and IoU (83.86) on the LEVIR-CD dataset.
I Introduction
Change detection in computer vision is a critical task that involves analyzing pairs of images captured at different times, known as bi-temporal images[1], to identify significant alterations in the scene. This process is vital for various applications, including monitoring urban development, environmental changes, and updating geographical information systems. The primary challenge in change detection lies in distinguishing meaningful changes, such as the construction of new infrastructure or alterations in land use, from variations caused by environmental factors like lighting conditions, weather, and seasonal changes, which introduce noise into the images [2].
Recent advancements have introduced innovative solutions to noise and its impact on feature extraction and change detection accuracy [3]. Among these, the state-of-the-art (SOTA) model DDPM-cd stands out because it incorporates a diffusion model for enhanced feature extraction. The Diffusion model is adept at filtering out the noise and highlighting the intrinsic characteristics of each image, creating a robust foundation for subsequent analysis. Once the features are extracted and refined through the Diffusion model, they are fed into a dedicated Change Detection Network. This network is designed to assess the feature sets from the bi-temporal images, determining the presence and extent of changes with a high degree of accuracy [4].
By leveraging the strengths of the Diffusion model for feature extraction [5], the DDPM-cd model effectively mitigates the influence of environmental noise, ensuring that the detected changes are indeed significant modifications within the scene rather than artifacts of varying capture conditions [6]. This approach enhances the reliability of change detection in complex environments, paving the way for more precise and efficient monitoring of changes across diverse landscapes.
II Prior Work
Denoising Diffusion Probabilistic Models (DDPMs) represent a sophisticated category within generative models, which, after adequate training, can produce images that closely mimic the training data distribution. These models distinguish themselves from generative model paradigms such as Generative Adversarial Networks (GANs), Variational Autoencoders (VAEs), autoregressive, and flow-based models by their training efficiency and versatility. DDPMs have seen extensive application across a wide range of image processing tasks, including but not limited to super-resolution, image deblurring, segmentation, colorization, inpainting, and semantic editing, as evidenced by their successful implementations [7, 8, 9, 10, 11].
Despite their proven efficacy in various domains of machine vision, the exploration of diffusion models within the context of remote sensing has been conspicuously absent. Given their remarkable success in mainstream computer vision tasks, diffusion models hold substantial promise for advancing the field of remote sensing, particularly in applications like change detection where identifying subtle changes over time is critical. This research aims to bridge this gap by demonstrating diffusion models’ applicability and potential benefits in remote sensing change detection. Our work highlights the versatility of diffusion models[12] and sets a precedent for their future exploration in remote sensing applications. In the ensuing sections, we will delve into the mechanics of the diffusion framework, setting the stage for a detailed discussion on its implementation and impact in remote sensing change detection.
Denoising diffusion probabilistic models represent a novel category within the broader spectrum of generative models, distinguished by their unique training mechanism and capability to produce images that closely mirror the distribution of the dataset they were trained on [13, 14, 15]. These models significantly depart from traditional generative approaches, such as Generative Adversarial Networks (GANs), Variational Autoencoders (VAEs), autoregressive models, and normalizing flows, by offering a more straightforward and efficient training process. Their versatility and effectiveness have led to widespread adoption across various image processing tasks, including but not limited to super-resolution, image deblurring, segmentation, colorization, inpainting, and semantic editing.
Despite the proven efficacy of diffusion models in various domains of machine vision, their application within remote sensing remains largely unexplored. This gap is notable given the potential benefits these models could bring to remote sensing, particularly in task change detection [16], where the ability to discern subtle differences between images over time is crucial. Our research aims to bridge this gap by demonstrating the applicability of diffusion models to change detection in remote sensing. By doing so, we expand the utility of diffusion models and introduce a promising new avenue for advancements in remote sensing technology. In the subsequent sections, we delve into the diffusion framework, laying the groundwork for understanding how these models can be effectively applied to remote sensing challenges.
Focusing on the operational mechanics, the denoising model, denoted as $M_{\delta}$, operates by taking a noisy image, ${u_{k}}$, as
$u_{k}=\sqrt{\xi_{k}}u_{0}+\sqrt{1-\xi_{k}}\delta$ (1)
where $\delta=\boldsymbol{N}(\boldsymbol{0},\boldsymbol{I})$. This initiates a process whereby the model incrementally reduces the noise level across a series of steps, effectively reversing the diffusion process. The gradual denoising procedure allows for reconstructing the original image from its noisy counterpart, showcasing the model’s capability to recover high-fidelity images from degraded inputs defined as
$\mathbb{D}_{u_{0},\delta}\left\|M_{\delta}(\tilde{u},k)-\delta\right\|_{2}^{2}$ (2)
The core functionality of denoising diffusion probabilistic models is pivotal for tasks that demand high levels of accuracy and detail in image restoration and analysis [17, 18], especially within the specialized field of remote sensing change detection  [19]. The inference process commences with an initial state of Gaussian noise, represented by $u_{k}$, and through iterative denoising across each timestep $k$, the model progressively recovers $u_{k-1}$ from $u_{k}$. This iterative refinement plays a critical role in reconstructing the original image from a state of significant degradation, underscoring the model’s effectiveness in applications where precision in image recovery is paramount as
$u_{k-1}\leftarrow{}\frac{1}{\sqrt{\alpha_{k}}}\left(u_{k}-\frac{1-\alpha_{k}}{\sqrt{1-\gamma_{k}}}M_{\delta}(u_{k},k)\right)+\xi_{k}\boldsymbol{q},$ (3)
where $\boldsymbol{q}=\boldsymbol{N}(\boldsymbol{0},\boldsymbol{I})$ and $k=T,\cdots,1$.
III Methodology
In examining the baseline model for change detection, we identified two significant limitations impacting its performance. Firstly, the model lacks a mechanism to integrate semantic information within individual temporal branches. Secondly, it fails to capture the temporal correlations between these branches. We introduced two novel feature attention mechanisms into the baseline architecture to mitigate these shortcomings. The enhancements comprise the Siamese Semantic Reasoning (Siam-SR) block [20] and the Cross-temporal Semantic Reasoning (Cot-SR) block [21] are shown on Fig.1a. The Siam-SR block is designed to process and encode semantic information specific to each temporal branch, ensuring that the model recognizes and utilizes the contextual details within individual time points. Concurrently, the Cot-SR block identifies and models the correlations across different temporal branches, facilitating a more cohesive understanding of the changes over time. Fig.2 and Fig.3 depict the architectural integration of these blocks, illustrating how they complement each other to enrich the model’s capacity for nuanced change detection.
While the Denoising Diffusion Probabilistic Model for Change Detection (DDPM-cd) achieves commendable efficacy, it ostensibly neglects to harness the intrinsic attributes of bi-temporal imagery. The methodology conventionally entails the assimilation of differential features derived from the diffusion model into the Change Detection Network (Classifier), conspicuously omitting a critical consideration for the ambient noise [22] introduced during the acquisition of the temporal images. This oversight precipitates a palpable need for a pre-emptive noise attenuation strategy to refine the input data quality.
In pursuit of this objective, we have elected to implement a strategy delineated by ChangerEx, as depicted in Figure 1b, named the Flow Dual-Alignment Fusion (FDAF). This sophisticated technique engenders the extraction of dual feature sets via the diffusion model [23], which are subsequently subjected to a convolutional network, specifically Flownet, followed by an image warping procedure before their integration into the Change Detection Network. The essence of this approach lies in its utilization of image warping to ameliorate the discrepancies between the paired images, thereby mitigating the impact of noise and enhancing the fidelity of the change detection process.
This advanced alignment fusion elevates the precision of feature representation and significantly diminishes the noise-induced anomalies, setting the stage for a more discerning and accurate change detection capability. Through the adoption of FDAF, we are poised to transcend the conventional limitations of DDPM-cd, imbuing the model with a refined sensitivity to the subtleties of bi-temporal image analysis [24] and fortifying its resilience against the vicissitudes of environmental conditions during image capture.
(a) Attention Based Model.
(b) FDAF Based Model
Figure 1: Structure of the baseline model with attention and FDAF.
IV Experiments
In our comprehensive investigation, we conducted an empirical analysis to ascertain the efficacy of change detection methodologies across two publicly available Remote Sensing datasets, LEVIR-CD and WHU-CD. Our evaluation metrics included the F1 score and the Intersection over Union (IoU) to quantitatively assess the accuracy in identifying altered regions within these datasets as shown in Fig.4. Based on Table-I, within the scope of the LEVIR-CD dataset, the foundational model yielded an F1 score of 90.91 and an IoU of 83.35. Enhancements introduced through the singular application of feature attraction augmented these metrics to 91.52 and 84.36, respectively. Conversely, the isolated deployment of the Flow Dual-Alignment Fusion (FDAF) precipitated a diminishment in performance, registering scores of 82.61 and 70.37. Notably, the concurrent application of both methodologies culminated in a marginal modification of the scores to 91.37 and 84.11, respectively.
Parallel observations were recorded with the WHU-CD dataset, where the baseline configuration delivered an F1 score of 92.65 and an IoU of 86.31. Applying feature attraction in isolation slightly reduced the metrics to 92.40 and 85.87, respectively. In stark contrast, the sole implementation of FDAF significantly eroded performance, with scores plummeting to 76.33 and 61.71. The amalgamation of feature attention and FDAF mechanisms resulted in adjusted scores of 90.27 and 82.27, respectively.
These findings illuminate the intricate dynamics of feature attention and FDAF within the context of change detection in remote sensing [25]. The augmentation of feature attention alone demonstrated a propitious impact on the LEVIR-CD dataset, underscoring its utility in enhancing model performance by intricately capturing and leveraging the temporal correlations between bi-temporal image sequences. Conversely, the integration of the FDAF network, while theoretically posited to ameliorate performance by mitigating noise through alignment and fusion mechanisms, empirically manifested a deleterious effect across both datasets. This dichotomy emphasizes the nuanced and dataset-specific efficacy of these advanced methodological interventions in remote sensing change detection.
Figure 2: Structure of Cot-SR blocks.
Figure 3: Structure of Siam-SR blocks.
LEVIR-CD WHU-CD
Method F1 IoU F1 IoU
88.44 82.17 91.24 84.70
90.18 83.86 91.74 84.93
81.57 69.26 71.82 57.60
Baseline Baseline+Attention Baseline+FDAF Baseline+Attention+FDAF 91.68 82.73 90.27 82.27
TABLE I: The average quantitative results on LEVIR-CD and WHU-CD. All the values are in percentage. Higher values of F1 and IoU indicate good CD performance.
(a) LEVIR, A.
(b) LEVIR, B.
(c) LEVIR, A+B.
(d) WHU, A.
(e) WHU, B.
(f) WHU, A+B.
Figure 4: Qualitative results. A is feature attention, B is FDAF.
To execute our experiments, a singular NVIDIA A6000 Ada Generation GPU was meticulously allocated to ensure uniform computational resources across all trials. The datasets utilized, specifically LEVIR-CD and WHU-CD, underwent preprocessing to standardize the input dimensions to 256x256 pixels, aligning with the established baseline methodology. Within the LEVIR-CD dataset, our training and validation cohorts comprised 7,120 and 2,048 instances, respectively, whereas the WHU-CD dataset was partitioned into 5,952 training and 1,488 validation samples.
To navigate the computational demands of our experiments, we adopted a batch size of eight during the training phase. The optimization of the neural networks was meticulously orchestrated utilizing the Binary Cross-Entropy (BCE) loss function, in conjunction with the AdamW optimizer, a decision underscored by its robustness in handling sparse gradients and regularization. Notably, the diffusion model’s parameters were preserved in a static state throughout the optimization process to isolate the effects of our experimental interventions.
The learning rate was strategically initialized at 1e-5, with a policy of linear decay implemented to gradually reduce this rate to zero as the training epochs progressed. This approach was designed to finely tune the model’s parameters without overshooting the optimal loss landscape. For the LEVIR dataset, the training regimen extended over 120 epochs, whereas for the WHU dataset, a slightly condensed schedule of 80 epochs was deemed sufficient. To ensure the capture of the most efficacious model configuration, we instituted a protocol to record the best validation score after each epoch, thereby enabling the identification and retention of the model iteration exhibiting the highest performance metrics.
V Conclusion
Within the ambit of this discourse, we elucidate a pioneering methodology for change detection, predicated on the nuanced manipulation of feature representations. Predominant paradigms have traditionally leveraged diffusion models as quintessential mechanisms for feature extraction, delineating changed regions within image datasets [26]. Our contemplation on the utility of feature maps has precipitated the inception of two novel augmentative networks: Feature Attention and Flow Dual-Alignment Fusion (FDAF), each tailored to refine the change detection process through strategic feature manipulation.
The Feature Attention mechanism is predicated on the hypothesis that discerning and learning the interrelations between dual feature maps can significantly amplify the efficacy of change detection [27]. This approach has manifested state-of-the-art (SOTA) performance across select datasets, underscoring the viability of correlational learning within feature attention frameworks [28]. Conversely, the FDAF paradigm was envisaged to mitigate environmental noise intrinsic to the conditions under which bi-temporal images are captured, thus enhancing the clarity of change detection. Despite its theoretical promise, empirical analyses have illuminated performance decrement issues with FDAF, ostensibly due to its propensity to filter out ambient noise and obfuscate pivotal feature distinctions between comparative images, thereby attenuating the precision of change detection.
In light of these findings, there lies an imperative for future investigation to refine the FDAF mechanism, ensuring that environmental noise can be effectively excised without compromising the integrity of target feature differences. Our research lays a foundational framework that, with iterative enhancements, could be expansively applied within the domain of change detection, offering robust methodologies for accurately identifying temporal and spatial alterations. The path forward involves ameliorating the identified limitations within the FDAF model and exploring our proposed methods’ adaptability and scalability across diverse change detection applications, heralding a new epoch in the evolution of change detection technologies.